# A Survey on Automatic Credibility Assessment of Textual Credibility Signals in the Era of Large Language Models

### v1

@@ 6.1. Datasets
The majority of the existing datasets were structured to address two related tasks: (i) classifying the input text as either containing persuasion techniques or not (binary task), and (ii) extracting all specific persuasion techniques present in the input text (multi-label task). The binary task offered a more straightforward approach, but it lacked the granularity necessary for a detailed analysis of propagandistic content. The multi-label task, on the other hand, was a more complex extension of the binary task, as it further required the model to identify the techniques individually, allowing a more fine-grained assessment of propaganda. For this reason, we focus our analysis on the multi-label task. A summary of existing datasets and their key attributes is presented in Table 6.
Dataset | # PTs | Language(s) | # Instances | Persuasion techniques | Content type | ||||||||||||||||||
SemEval-2019 (Da San Martino et al., 2019a) | 18 | English | 7,485 |
News articles | ||||||||||||||||||
Baisa et al. (2019) | 18 | Czech | 7,494 |
News articles | ||||||||||||||||||
@@
Lawson et al. (2020) | 4 | English | 90 | Authority, Commitment/Consistency, Liking, Scarcity | Emails | ||||||||||||||||||
PTC (SemEval-2020) (Da San Martino et al., 2020a) | 14 | English | 8,981 |
News articles | ||||||||||||||||||
SemEval-2021 (Dimitrov et al., 2021) | 22 | English | 2,488 |
Facebook posts | ||||||||||||||||||
WANLP-2022 (Alam et al., 2022b) | 20 | Arabic | 1,942 |
Tweets | ||||||||||||||||||
Macagno (2022) | 9 |
2,657 |
Tweets | ||||||||||||||||||
ArAIEval-2023 (Hasanain et al., 2023) | 23 | Arabic | 5,919 |
SemEval-2023 (Piskorski et al., 2023) | 23 |
49,444 |
News articles | ||||||||||||||||||
Almotairy et al. (2024) | 15 | Arabic | 2,100 |
Tweets |
@@
Macagno (2022) introduced the first multilingual dataset for detection of persuasion techniques, containing tweets in English, Italian, and Portuguese. However, the dataset is small ( instances in total) for the purpose of training deep learning models, and similarly to (Baisa et al., 2019), the set of persuasion techniques considered differs substantially from other datasets. SemEval-2023 (Piskorski et al., 2023) introduced a large-scale multilingual dataset covering languages (English, French, German, Georgian, Greek, Italian, Polish, Russian, and Spanish), with almost news articles labelled with different persuasion techniques, therefore being the largest dataset currently available in terms of number of instances, languages, and persuasion techniques. Three languages were considered ”surprise languages” (Georgian, Greek, and Spanish), for which training sets were not available during the competition, thus encouraging systems to deal with out-of-domain data. Furthermore, the set of persuasion techniques were grouped into coarse-grained categories. For example, the techniques of Loaded Language, Obfuscation, Intentional Vagueness, Confusion, Exaggeration or Minimisation, and Repetition, were grouped into the umbrella of Manipulative Wording.
@@ 6.2. Methods and models
The majority of methods and models developed for the task of automatic detection of persuasion techniques were introduced in the shared tasks discussed in the previous section. The joint effort of multiple different attempts at producing the best-scoring system allows to identify which methodological decisions are key to producing accurate models to detect persuasion. Table 7 summarises the highest-scoring202020Shared task systems that did not publish a description of their approach are not considered in our analysis. systems across the shared tasks aimed at automatic detection of persuasion techniques.
System | Dataset | Model | Approach | F1 Micro | |||||||||||||
newspeak (Yoosuf and Yang, 2019) | SemEval-2019 (Da San Martino et al., 2019b) | BERT base uncased |
@@
Hitachi (Morio et al., 2020) | SemEval-2020 (Da San Martino et al., 2020a) |
@@
MinD (Tian et al., 2021) | SemEval-2021 (Dimitrov et al., 2021) |
@@
NGU CNLP (Hussein et al., 2022) | WANLP-2022 (Alam et al., 2022b) | AraBERT |
@@
UL & UM6P (Lamsiyah et al., 2023) | ArAIEval-2022 (Hasanain et al., 2023) | AraBERT-Twitter-v2 |
@@
Razuvayevskaya et al. (2024) | SemEval-2023 (Piskorski et al., 2023) | XLM-RoBERTa large |
@@
Arguably the most relevant methods employed in top-performing systems are aimed towards dealing with data skewness. Most datasets for this task suffer from data skewness, meaning the majority of persuasion techniques are underrepresented, while a small set of persuasion techniques comprise the majority of the dataset. To deal with this issue, several different methods are employed, such as oversampling underrepresented techniques (Yoosuf and Yang, 2019; Ek and Ghanimifard, 2019; Pauli et al., 2023), scaling the contribution of the techniques to the loss function according to their proportion (i.e., class weighting) (Yoosuf and Yang, 2019; Morio et al., 2020; Abdel-Salam, 2023; Razuvayevskaya et al., 2024), and using supplementary data obtained with (i) data augmentation (Gupta and Sharma, 2021), (ii) semi-supervised methods (Jurkiewicz et al., 2020), or (iii) similar datasets (Tian et al., 2021; Abujaber et al., 2021; Hussein et al., 2022; Mittal and Nakov, 2022).
@@ 6.3. Tools and services
The Propaganda Persuasion Techniques Analyzer (PRTA) (Da San Martino et al., 2020b) was designed to detect instances of propaganda in texts by highlighting the spans where specific techniques were used. It provided users with the ability to compare texts based on their use of propaganda techniques, offering detailed statistics on the prevalence of these techniques, both overall and over time. The tool allowed for filtering based on time intervals, keywords, or political orientation of the media. Users could analyse texts either through a dedicated interface or via an API212121https://www.tanbih.org/prta. PRTA used a BERT model trained on the SemEval-2019 dataset, fine-tuned for both fragment-level and sentence-level classification tasks. To gather data, PRTA crawled a growing list of over RSS feeds, Twitter accounts, and websites, extracting text via the Newspaper3k library and performing deduplication using a hash function. The system then identified sentences containing propaganda, organised the articles into topics (such as COVID-19 or Brexit), and allowed users to compare the use of propaganda techniques across various media sources.
The GATE Cloud infrastructure (Cunningham et al., 2002) operated by the University of Sheffield provides a service to detect different persuasion techniques across multiple languages using a multilingual BERT (mBERT) model222222https://cloud.gate.ac.uk/shopfront/displayItem/persuasion-classifier. The model was fine-tuned using data from SemEval-2023 in English, French, German, Italian, Polish, and Russian. Class weighting was applied during fine-tuning to account for the label skewness issue. The tool was capable of processing up to documents per day free of charge through its API, with an average processing rate of 2 documents per second. Researchers could request higher quotas if needed for larger-scale analysis.
@@ 6.4. Discussion
Most commonly adopted persuasion techniques. Table 6 shows that out of the datasets present a similar set of persuasion techniques (Da San Martino et al., 2019b, 2020a; Dimitrov et al., 2021; Alam et al., 2022b; Hasanain et al., 2023; Piskorski et al., 2023; Almotairy et al., 2024), while Baisa et al. (2019), Lawson et al. (2020), and Macagno (2022) largely diverge from them. Considering these seven similar datasets, we observe a set of persuasion techniques that are consistent among them: Appeal to authority, Appeal to fear/prejudice, Causal oversimplification, Doubt, Exaggeration or minimization, Flag-waving, Loaded language, Name calling or labeling, Slogans, and Whataboutism appear in all seven datasets, while Repetition appears in six datasets (except Almotairy et al. (2024)), and Reductio ad Hitlerum appears in five datasets (except Hasanain et al. (2023), and Piskorski et al. (2023)). Combining existing datasets with overlapping persuasion techniques could be a promising research direction to develop more robust models and benchmarking resources. For example, Tian et al. (2021) achieved 1st place in SemEval-2021 Task 6 by leveraging the dataset from Da San Martino et al. (2019b) as additional training data.
@@
Label skewness. A common characteristic inherent to this label scheme is the significant skewness in distribution of the persuasion techniques. For example, in Piskorski et al. (2023) (the largest multilingual dataset, with instances), a small set of techniques represent of the entire dataset - Loaded Language (), Name Calling-Labelling (), Casting Doubt (), Questioning the Reputation (), Appeal to Fear-Prejudice (), and Exageration/Minimisation () - while the remaining techniques account for only of the dataset. The same trend can be seen in Da San Martino et al. (2019b), with techniques representing of the dataset - Loaded language (), Name calling, labeling (), Repetition (), Exaggeration, minimization (), Doubt (), and Appeal to fear/prejudice () - while the remaining techniques account for of the dataset. This sparse label distribution makes training and evaluating models more challenging, as the majority of persuasion techniques are considerably underrepresented.
@@
Evaluation. Shared tasks on persuasion techniques typically employ the F1-Micro measure as the official evaluation metric (Dimitrov et al., 2021; Da San Martino et al., 2020a; Alam et al., 2022b; Hasanain et al., 2023; Piskorski et al., 2023), however, the F1-Micro does not account for skewed class distributions, as opposed to F1-Macro232323Generally F1-Macro is also reported as a secondary metric, but for the purpose of the competition, the F1-Micro determines which system is best.. This evaluation strategy encourages models that excel at predicting a few densely distributed persuasion techniques. For example, in ArAIEval-2023 (Hasanain et al., 2023), the baseline model (majority vote classifier that always predicts the most common persuasion technique), achieves an F1-Micro of , and an F1-Macro of , representing a difference of between the two scores. Similarly, the best submission by Lamsiyah et al. (2023) achieves F1-Micro and F1-Macro scores of and , a difference of . Similar figures are seen in other shared tasks such as SemEval-2023 (Piskorski et al., 2023), with the average F1-Micro score for first place submissions across all languages resulting in , and the F1-Macro in , a difference of . These figures highlight how current state-of-the-art models trained for this task still lack the ability to predict underrepresented persuasion techniques effectively.
@@
7. Claims and veracity
Given the impossibility to verify the credibility of every single piece of information posted online, two credibility signals and corresponding tasks are particularly crucial in supporting fact-checkers and other players involved in the fight against disinformation and misinformation (Guo et al., 2022): check-worthiness detection and the retrieval of previously fact-checked claims. The former is aimed at identifying which claims are check-worthy because of their relevance and interest to the general public (Hassan et al., 2015). The second, which ideally takes place after the first one, is meant to ease claim verification by checking whether the given claim was already verified before by searching in existing repositories of fact-checked claims (Shaar et al., 2020), both monolingual and multilingual.
Outputs of both tasks can also be used (usually in combination) as credibility signals to indicate whether the check-worthy (central) claims contained in a piece of content have been fact-checked and if so, what was the given veracity value or values. They are recognized as such both in the credibility signals list created by the Credible Web Community Group (signals ‘article has a central claim’ and ‘fact-check status of a claim’; see (Multiple authors and Hawke, 2019)) as well as in the related works (Zhang et al., 2018). The check-worthiness of a claim can be considered a content-based signal, while fact-check status of a claim is a context-based one, since it requires additional external sources to be detected.

### v2

@@ 6.1. Datasets
The majority of existing datasets are designed to support two related tasks: (i) classifying the input text as either containing persuasion techniques or not (binary classification), and (ii) extracting all specific persuasion techniques present in the input text (multi-label classification). While the binary task provides a simpler framework, it lacks the necessary granularity for detailed analysis of propagandistic content. The multi-label task is a more complex extension, requiring the model to identify individual persuasion techniques and thus enabling a more nuanced evaluation. Importantly, these tasks are implemented at different levels of granularity across datasets – some annotate propaganda at the token-level (fine-grained), others at the sentence-level (coarse-grained), and some at the fragment-level (mid-grained), where continuous spans of text within a sentence are labelled. These distinctions significantly affect how models are trained and evaluated. For this reason, and due to its relevance for fine-grained analysis, we focus our analysis on the multi-label task. A summary of existing datasets and their key attributes, including annotation granularity, is presented in Table 6.
| Dataset | Language(s) | # Instances | Content type | Classes | |||||||||
| SemEval-2019 (Da San Martino et al., 2019a) | English | 7,485 | News articles | 18 persuasion techniques: Appeal to authority, Appeal to fear/prejudice, Bandwagon, Black-and-white fallacy/dictatorship, Causal oversimplification, Doubt, Exaggeration or minimization, Flag-waving, Loaded language, Name calling or labelling, Obfuscation/intentional vagueness/confusion, Red herring, Reductio ad Hitlerum, Repetition, Slogans, Straw man, Thought-terminating cliché, Whataboutism | |||||||||
| Baisa et al. (2019) | Czech | 7,494 | News articles | 18 persuasion techniques: Blaming, Labelling, Argumentation, Emotions, Demonizing, Relativizing, Fear mongering, Fabulation, Opinion, Location, Source, Russia, Expert, Attitude to a politician, Topic, Genre, Focus, Overall sentiment | |||||||||
@@
| Lawson et al. (2020) | English | 90 | Emails | 4 persuasion techniques: Authority, Commitment/Consistency, Liking, Scarcity | |||||||||
| PTC (SemEval-2020) (Da San Martino et al., 2020a) | English | 8,981 | News articles | 14 persuasion techniques: Appeal to authority, Appeal to fear/prejudice, Bandwagon/Reductio ad Hitlerum, Black-and-white fallacy, Causal oversimplification, Doubt, Exaggeration/minimization, Flag-waving, Loaded language, Name calling/labelling, Repetition, Slogans, Thought-terminating cliché, Whataboutism/Straw man/Red herring | |||||||||
| SemEval-2021 (Dimitrov et al., 2021) | English | 2,488 | Facebook posts | 22 persuasion techniques: Appeal to authority, Appeal to (Strong) Emotions, Appeal to Fear or Prejudices, Bandwagon, Black-and-White Fallacy or Dictatorship, Causal Oversimplification, Flag-Waving Doubt, Exaggeration or Minimisation, Slogans, Glittering Generalities (Virtue), Loaded Language, Misrepresentation of Someone’s Position (Straw Man), Name Calling or Labelling, Repetition, Obfuscation/Intentional Vagueness/Confusion, Presenting Irrelevant Data (Red Herring), Reductio ad Hitlerum, Thought-Terminating Cliché, Smears, Transfer, Whataboutism | |||||||||
| WANLP-2022 (Alam et al., 2022b) | Arabic | 1,942 | Tweets | 20 persuasion techniques: Appeal to authority, Appeal to fear/prejudices, Bandwagon, Black-and-white fallacy/dictatorship, Flag-waving, Doubt Causal oversimplification, Exaggeration/minimisation, Glittering generalities (virtue), Loaded language, Misrepresentation of someone’s position (straw man), Name calling or labelling, Repetition, Slogans Obfuscation/intentional vagueness/confusion, Presenting irrelevant data (red herring), Reductio ad Hitlerum, Smears, Thought-terminating cliché, Whataboutism | |||||||||
| Macagno (2022) |
2,657 | Tweets | 9 persuasion techniques: Straw man, False dichotomy, Ignoring qualifications, Question begging epithets, Post hoc ergo propter hoc, Hasty generalization, Slippery Slope, Persuasive definition, Quasi-definition | |||||||||
| ArAIEval-2023 (Hasanain et al., 2023) | Arabic | 5,919 |
23 persuasion techniques: Appeal to Authority, Appeal to Fear/Prejudice, Appeal to Hypocrisy, Appeal to Popularity, Appeal to Time, Appeal to Values, Casting Doubt, Causal Oversimplification, Consequential Oversimplification, Conversation Killer, Exaggeration or Minimisation, False Dilemma or No Choice, Flag Waving, Guilt by Association, Loaded Language, Name Calling or Labelling, Red Herring, Repetition Obfuscation/intentional vagueness/confusion, Slogans Questioning the Reputation, Strawman, Whataboutism | |||||||||
| SemEval-2023 (Piskorski et al., 2023) |
49,444 | News articles | 23 persuasion techniques: Appeal to Authority, Appeal to Fear/Prejudice, Appeal to Hypocrisy, Appeal to Popularity, Appeal to Time, Appeal to Values, Casting Doubt, Causal Oversimplification, Consequential Oversimplification, Conversation Killer, Exaggeration or Minimisation, False Dilemma or No Choice, Flag Waving, Guilt by Association, Loaded Language, Name Calling or Labelling, Red Herring, Repetition Obfuscation/intentional vagueness/confusion, Questioning the Reputation, Slogans, Strawman, Whataboutism | |||||||||
| Almotairy et al. (2024) | Arabic | 2,100 | Tweets | 15 persuasion techniques: Flag-waving, Smears, Name-calling, Loaded language, Exaggeration, Whataboutism, Glittering, Doubt, Causal oversimplification, Dictatorship, Appeal to fear, Slogan, Thought-terminating cliché, Appeal to authority, Reductio ad Hitlerum |
@@
Macagno (2022) introduced the first multilingual dataset for detection of persuasion techniques, containing tweets in English, Italian, and Portuguese. However, the dataset is small ( instances in total) for the purpose of training deep learning models, and similarly to (Baisa et al., 2019), the set of persuasion techniques considered differs substantially from other datasets. SemEval-2023 (Piskorski et al., 2023) introduced a large-scale multilingual dataset covering languages (English, French, German, Georgian, Greek, Italian, Polish, Russian, and Spanish), with almost news articles labelled with different persuasion techniques, therefore being the largest dataset currently available in terms of number of instances, languages, and persuasion techniques. Three languages were considered “surprise languages” (Georgian, Greek, and Spanish), for which training sets were not available during the competition, thus encouraging systems to deal with out-of-domain data. Furthermore, the set of persuasion techniques were grouped into coarse-grained categories. For example, the techniques of Loaded Language, Obfuscation, Intentional Vagueness, Confusion, Exaggeration or Minimisation, and Repetition, were grouped into the umbrella of Manipulative Wording.
@@ 6.2. Methods and models
The majority of methods and models developed for the task of automatic detection of persuasion techniques were introduced in the shared tasks discussed in the previous section. The joint effort of multiple different attempts at producing the best-scoring system allows to identify which methodological decisions are key to producing accurate models to detect persuasion. Table 7 summarises the highest-scoring191919Shared task systems that did not publish a description of their approach are not considered in our analysis. systems across the shared tasks aimed at automatic detection of persuasion techniques.
| Dataset | System | Model | Approach | F1 Micro | |||||||||||||
| SemEval-2019 (Da San Martino et al., 2019b) | newspeak (Yoosuf and Yang, 2019) | BERT base uncased |
@@
| SemEval-2020 (Da San Martino et al., 2020a) | Hitachi (Morio et al., 2020) |
@@
| SemEval-2021 (Dimitrov et al., 2021) | MinD (Tian et al., 2021) |
@@
| WANLP-2022 (Alam et al., 2022b) | NGU CNLP (Hussein et al., 2022) | AraBERT |
@@
| ArAIEval-2022 (Hasanain et al., 2023) | UL & UM6P (Lamsiyah et al., 2023) | AraBERT-Twitter-v2 |
@@
| SemEval-2023 (Piskorski et al., 2023) | Razuvayevskaya et al. (2024) | XLM-RoBERTa large |
@@
Arguably the most relevant methods employed in top-performing systems are aimed towards dealing with data skewness. Most datasets for this task suffer from class imbalance, meaning the majority of persuasion techniques are underrepresented, while a small set of persuasion techniques comprise the majority of the dataset. To deal with this issue, several different methods are employed, such as oversampling underrepresented techniques (Yoosuf and Yang, 2019; Ek and Ghanimifard, 2019; Pauli et al., 2023), scaling the contribution of the techniques to the loss function according to their proportion (i.e., class weighting) (Yoosuf and Yang, 2019; Morio et al., 2020; Abdel-Salam, 2023; Razuvayevskaya et al., 2024), and using supplementary data obtained with (i) data augmentation (Gupta and Sharma, 2021), (ii) semi-supervised methods (Jurkiewicz et al., 2020), or (iii) similar datasets (Tian et al., 2021; Abujaber et al., 2021; Hussein et al., 2022; Mittal and Nakov, 2022).
@@ 6.3. Tools and services
The Propaganda Persuasion Techniques Analyzer (PRTA) (Da San Martino et al., 2020b) is a discontinued tool that is no longer available online. It was designed to detect instances of propaganda in texts by highlighting the spans where specific techniques were used. It provided users with the ability to compare texts based on their use of propaganda techniques, offering detailed statistics on the prevalence of these techniques, both overall and over time. PRTA used a BERT model trained on the SemEval-2019 dataset, fine-tuned for both fragment-level and sentence-level classification tasks. To gather data, PRTA crawled a growing list of over RSS feeds, Twitter accounts, and websites, extracting text via the Newspaper3k library and performing deduplication using a hash function. The system then identified sentences containing propaganda, organised the articles into topics (such as COVID-19 or Brexit), and allowed users to compare the use of propaganda techniques across various media sources. The tool also allowed for filtering based on time intervals, keywords, or political orientation of the media. Although the user interface of the PRTA tool is no longer available, the underlying API for propaganda detection (along with other credibility signals) remains accessible202020https://apihub.tanbih.org/propaganda/.
The GATE Cloud infrastructure (Cunningham et al., 2002) operated by the University of Sheffield provides a service to detect different persuasion techniques across multiple languages using a multilingual BERT (mBERT) model212121https://cloud.gate.ac.uk/shopfront/displayItem/persuasion-classifier. The model was fine-tuned using data from SemEval-2023 in English, French, German, Italian, Polish, and Russian. Class weighting was applied during fine-tuning to account for the label skewness issue. The tool is capable of processing up to documents per day free of charge through its API, with an average processing rate of 2 documents per second. Researchers can request higher quotas if needed for larger-scale analysis.
@@ 6.4. Discussion
Most commonly adopted persuasion techniques. Table 6 shows that out of the datasets present a similar set of persuasion techniques (Da San Martino et al., 2019b, 2020a; Dimitrov et al., 2021; Alam et al., 2022b; Hasanain et al., 2023; Piskorski et al., 2023; Almotairy et al., 2024), while (Baisa et al., 2019; Lawson et al., 2020; Macagno, 2022) largely diverge from them. Considering these seven similar datasets, we observe a set of persuasion techniques that are consistent among them: Appeal to authority, Appeal to fear/prejudice, Causal oversimplification, Doubt, Exaggeration or minimization, Flag-waving, Loaded language, Name calling or labelling, Slogans, and Whataboutism appear in all seven datasets, while Repetition appears in six datasets (except (Almotairy et al., 2024)), and Reductio ad Hitlerum appears in five datasets (except (Hasanain et al., 2023; Piskorski et al., 2023)). Combining existing datasets with overlapping persuasion techniques could be a promising research direction to develop more robust models and benchmarking resources. For example, Tian et al. (2021) achieved 1st place in SemEval-2021 Task 6 by leveraging the dataset from (Da San Martino et al., 2019b) as additional training data.
@@
Label skewness. A common characteristic inherent to this label scheme is the significant skewness in distribution of the persuasion techniques. For example, in SemEval-2023 (Piskorski et al., 2023) (the largest multilingual dataset, with instances), a small set of techniques represent of the entire dataset - Loaded Language (), Name Calling-Labelling (), Casting Doubt (), Questioning the Reputation (), Appeal to Fear-Prejudice (), and Exageration/Minimisation () - while the remaining techniques account for only of the dataset. The same trend can be seen in (Da San Martino et al., 2019b), with techniques representing of the dataset - Loaded language (), Name calling, labelling (), Repetition (), Exaggeration, minimization (), Doubt (), and Appeal to fear/prejudice () - while the remaining techniques account for of the dataset. This sparse label distribution poses several challenges for training and evaluation. In particular, the overrepresentation of a few dominant techniques may lead models to bias their predictions toward these frequent categories, while underrepresented techniques suffer from insufficient training examples. This data imbalance increases the risk of overfitting to the few available samples of rare classes, potentially resulting in poor generalization.
@@
Evaluation. Shared tasks on persuasion techniques detection typically employ the F1-Micro measure as the official evaluation metric (Dimitrov et al., 2021; Da San Martino et al., 2020a; Alam et al., 2022b; Hasanain et al., 2023; Piskorski et al., 2023), however, the F1-Micro does not account for skewed class distributions, as opposed to F1-Macro222222Generally F1-Macro is also reported as a secondary metric, but for the purpose of the competition, the F1-Micro determines which system is best.. This evaluation strategy encourages models that excel at predicting a few densely distributed persuasion techniques. For example, in ArAIEval-2023 (Hasanain et al., 2023), the baseline model (majority vote classifier that always predicts the most common persuasion technique), achieves an F1-Micro of , and an F1-Macro of , representing a difference of between the two scores. Similarly, the best submission by Lamsiyah et al. (2023) achieves F1-Micro and F1-Macro scores of and , a difference of . Similar figures are seen in other shared tasks such as SemEval-2023 (Piskorski et al., 2023), with the average F1-Micro score for first place submissions across all languages resulting in , and the F1-Macro in , a difference of . These figures highlight how current state-of-the-art models trained for this task still lack the ability to predict underrepresented persuasion techniques effectively.
@@
Adoption of LLMs. Recent research has increasingly explored the capabilities of large language models in generating persuasive and deceptive content, particularly in the context of disinformation and propaganda (Simchon et al., 2024; Breum et al., 2024; Rogiers et al., 2024; Xu et al., 2024; Pauli et al., 2024). LLMs can manipulate existing factual content into misleading or false narratives at scale, often outperforming human-written disinformation in believability and fluency (Matz et al., 2024; Chen and Shu, 2023) and can even personalize the content to specific target groups (Zugecova et al., 2025). However, while the generation side has been widely studied, significantly less attention has been paid to using LLMs for detecting or countering persuasive techniques. Bridging this gap by harnessing LLMs for persuasion detection presents a promising but underexplored research direction.
7. Check-worthy and fact-checked claims
Given the impossibility to verify the veracity of every single piece of information posted online, two credibility signals and corresponding tasks are particularly crucial in supporting fact-checkers and other actors involved in the fight against disinformation and misinformation (Guo et al., 2022): check-worthiness detection and the retrieval of previously fact-checked claims. The former is aimed at identifying which claims are check-worthy because of their relevance and interest to the general public (Hassan et al., 2015). The second, which ideally takes place after the first one, is meant to ease claim verification by checking whether the given claim was already verified before by searching in existing repositories of fact-checked claims (Shaar et al., 2020), both monolingual and multilingual.
While these tasks are usually studied in separation as a part of the fact-checking pipelines, their outputs can be and also are used (usually in combination) as credibility signals to indicate whether the check-worthy (central) claims contained in a piece of content have been fact-checked and if so, what was the given veracity value or values. In fact, they are recognized as such both in the credibility signals list created by the Credible Web Community Group (signals ‘article has a central claim’ and ‘fact-check status of a claim’; see (Multiple authors and Hawke, 2019)) as well as in the related works (Zhang et al., 2018). The check-worthiness of a claim can be considered a content-based signal, while fact-check status of a claim is a context-based one, since it requires additional external sources to be detected. On the other hand, fact-checking itself is not considered a credibility signal per se – its aim is to directly ascertain the veracity value of a piece of information, while the former are only signals that the information might be check-worthy (and thus more attention is needed before the users make credibility judgment) or that it contains a claim that have already been fact-checked (but it does not by itself analyse the stance towards that claim, only its presence).
It is also worth noting that these signals differ from factuality signal as it is understood in the surveyed works as discussed in Section 5).